OpenStudy (anonymous):

joe gives out 22 newspapers everyday in september and october how many papers did he give out in all? joe gives out 22 newspapers everyday in september and october how many papers did he give out in all? @Mathematics

OpenStudy (anonymous):

22x 30 + 22x 31 = 22 (30+31) = 22 (61) =1342